OpenStudy (anonymous):

y varies jointly with p and q. When p = 12 and q = 4, y = 144. Find y when p = 10 and q = 6

OpenStudy (anonymous):

y=p^q/2=10^3=1000

OpenStudy (anonymous):

y= 3(p X q) = 144 so, when p=10 & q=6 then y= 3(10*6) equals 180; that's what I'm thinking.

OpenStudy (anonymous):

there are many solution so i make a solution y=f(p,q)